Find the power series for f(z) = (1/(z^2)) around the point zo = i and indicate the radius of convergence
Answer #1

-| マーシャ2)2 (2-iP+ 2202-3)+2 一 it anan f(2) of 2 のりで i it an-ni 2020/7/2921:16 (2-2P+22(2ーモ) ー(レーし、マシュでくること) t =-1-22(2-i)+302

I use (1-x) ^-1= 1+x+x^2+x^3+.........